Suppose the economy is described by following characteristics
C(Y)=200+0.4(Y-T)
I=100
G=400
T=100
Determine the equilibrium output level.
1
Expert's answer
2020-02-10T08:48:45-0500

Y - the equilibrium output level

Y = c0 + c1(Y − T) + G + I

c0 = 200

c1 = 0.4

Y = 200 + 0.4(Y - 100) + 400 + 100

Y = 200 + 0.4Y - 40 + 400 + 100

0.6Y = 660

Y = 1100

Answer: Y=1100